Write the equation of the line passing through -3,-5 and 3,0.
Answer by Nate(3500)

Write the equation of the line passing through -3,-5 and 3,0.
M = (y - y)/(x - x) = (0 + 5)/(3 + 3) = (5)/(6)
y - 0 = (5/6)(x - 3)
y = (5/6)x - 2.5